How do you solve for each variable:
(x-3)(x-7)=0
b^2+3b-4=0
n^2+n-12=0
5x^2+18x=8
3a^2+4a=2a^2-2a-9

Answer by checkley77(12844)

(x-3)(x-7)=0
x-3=0
x=3 answer
x-7=0
x=7 answer.
-------------------------------------
b^2+3b-4=0
(b+4)(b-1)=0
b+4=0
b=-4 answer.
b-1=0
b=1 answer.
------------------------------------------------
Here's 2 solutions.
I'll leave the rest of your homework for you to practice on.